# Imaging of I Zw 18 by JWST: I. Strategy and First Results of Dusty Stellar Populations

### v1

@@ 4.1 Images
Our NIRCam-only image (Figure 2) reveals an extreme population of bright, blue, recently-formed massive stars, with positions centered upon two primary lobes of star formation, consistent with earlier optical/near-IR studies of I Zw 18 utilizing HST (e.g., Hunter & Thronson, 1995; Östlin, 2000; Cannon et al., 2002; Izotov & Thuan, 2004). An assortment of vivid red stellar sources distributed about the galaxy accentuates the presence of luminous evolved stars. Also apparent are the large supershells of dust and ionized gas carved out by previous star-formation and SNe activity. These align with extant maps of emission, including continuum-subtracted H and H from Cannon et al. (2002), which found significant concentrations of dust along such structures. Resolved background galaxies are apparent throughout the image spanning a variety of orientations and morphology.
@@
4.2 Color-Magnitude Diagrams
@@
We construct near-IR CMDs using NIRCam data which enabled our investigation of the various stellar populations inhabiting I Zw 18 (Figure 7). Our primary diagnostic CMD is F200W vs. F115W – F200W, which possesses the highest sensitivity to faint sources and therefore offers the largest potential pool of candidate stars ( = 689). This combination of filters was employed as a JWST-specific facsimile for standard diagnostic plots, enabling comparison with previous dusty and evolved star studies utilizing Spitzer bands such as SAGE and DUSTiNGS (e.g., Blum et al., 2006; Boyer et al., 2011, 2015a, 2015b). In addition, the methods executed by this study mirror recent stellar populations imaging work undertaken with JWST of other extragalactic sources, including NGC 6822 (Lenkić et al., 2023; Nally et al., 2023) and NGC 346 (Jones et al. 2023; Habel et al. submitted).
@@
The structure of points populating this CMD can be roughly described as two primarily vertical columns extending from a common foundation, which terminates as the detection threshold is reached at magnitudes fainter than F200W 28.0. At this level, the sensitivity limitations of the photometry results in a broad scatter. Points which make up the blueward (left-side) column of sources represent young stars of the Upper Main Sequence (UMS), with those exhibiting the brightest magnitudes being the most massive. The redward column of sources is principally populated by evolved stars. Here, fainter sources are on the Red Giant Branch (RGB), while at brighter luminosities these stars are chiefly He core burning RSGs and those on the AGB. Based on the distance modulus to I Zw 18, the Tip of the RGB (TRGB) is estimated to be located at F200W 25.6. A luminosity function for all galaxy stars (gray) and only those stellar sources populating the evolved star column of sources in this CMD (red) is presented as Figure 8. An apparent discontinuity of sources at F200W 25.6 supports this estimation of the TRGB location, though a combination of large photometric uncertainties and small-number statistics prevents definitive identification of the TRGB.
@@
In order to assist our preliminary point-source type classifications, we have included for reference parsec isochrones of an XMP stellar population ( = 0.0003; e.g., Bressan et al., 2012) at ages of 1 Myr, 10 Myr, 25 Myr, 50 Myr, 100 Myr, 250 Myr, 500 Myr, 1 Gyr, 2 Gyr, 5 Gyr, and 10 Gyr, overlaid on this CMD (see Figure 9). The distribution of sources seen reflects the existence of both older and younger populations: For a thorough examination of the star formation history (SFH) of this galaxy as discerned by the photometric data obtained with NIRCam, see Bortolini et al. (in prep.). The overall numbers for point-source identifications described in the following subsections from the young star and evolved star branches, including their respective dust-enshrouded candidates (§4.2.1), plus YSO and YSO cluster candidates (§4.2.2), are summarized in Table 4.
@@ 4.2.1 Dusty Evolved Star Candidates
Identification of dust-enshrouded stellar source candidates in I Zw 18 using IR photometry was accomplished by examining a CMD employing our most widely-separated NIRCam filter data (F444W vs. F115W – F444W; Figure 10). At these longer wavelengths the numbers of detected point sources ( = 183) diminish in comparison to our primary diagnostic CMD. This large color baseline, however, affords a robust separation between dusty and non-dusty stellar types: Any source detection with a color value F115W – F444W 0.0 may be considered to be at least somewhat enshrouded in dust. In order to account for photometric uncertainties, multiple ages, variance in metallicities, and other broadening effects, we have elected to implement a more conservative color cut at F115W – F444W = 0.5.
@@
For our list of dust-enshrouded source candidates identified from this CMD, we additionally required each detection to have been found in the F200W vs. F115W – F200W CMD located above the TRGB (F200W = 25.6). We note that, while some additional bona fide dust-enshrouded source candidates are likely to lie below this TRGB, we have elected to exclude them here for the sake of remaining conservative with our estimates, and to reflect larger photometric uncertainties at these magnitudes. For objects inhabiting the right-side (evolved stars) branch of the F200W vs. F115W – F200W CMD, those which are additionally identified as being dusty in the F444W vs. F115W – F444W CMD are expected to be AGB stars, RSGs, and bright YSO clusters. The fewer dusty sources of the left-side (young stars) branch of the F200W vs. F115W – F200W CMD are likely RSGs. In total, we find = 119 dusty source candidates in the F444W vs. F115W – F444W CMD, which are signified both in Figures 7 and 10 as overlaid green dots. Of these, = 99 belong to the evolved star branch, while = 20 belong to the young star branch.
@@
Stars inhabiting the AGB represent the final evolutionary stage of low- to intermediate-mass (0.6–10 M) MS progenitors, and are believed to be a major contributor of interstellar dust. Of interest to this observing program in particular are thermally-pulsing (TP-AGB) stars which lie above the TRGB. We may classify these AGB stars into two major categories: Oxygen- and carbon-rich (O-rich and C-rich, respectively). While O-rich AGB stars (5–10 M; 30–50 Myr) are capable of producing dust relatively quickly after formation, they are not expected to seed dust formation in metal-poor environments like the slower-evolving, lower-mass C-rich AGB stars (1.5–5 M; 200 Myr–3.6 Gyr), which create their own carbon (Boyer et al., 2015a). O- and C-rich TP-AGB stars may be distinguished from one other based on the dominant chemistries of their photospheres. If the value of the ratio of carbon to oxygen (C/O) is 1, the AGB star is C-rich, while a C/O value 1 is O-rich. Dust species borne from these populations depend upon the dominant element, such that carbonaceous dust grains are produced by C-rich AGB stars, while silicates are produced by O-rich AGB stars.
@@
Without the wider color baselines afforded by MIRI photometry, robust segregation between O- and C-rich AGB stars is not possible. Many of the evolved-star branch sources identified in the F200W vs. F115W – F200W CMD (Figure 7; red points), however, are expected to be TP-AGB stars. In addition, considering the extremely low metallicity of I Zw 18, many of these TP-AGB candidate stars are expected to be C-rich. The distribution of dust-enshrouded evolved stars (red points with over-plotted green dots) permeates the parameter space occupied by the F200W vs. F115W – F200W CMD’s evolved-star branch, suggesting the presence of many dusty C-rich AGB star candidates in I Zw 18.
@@
The locations of the evolved-star branch sources in this galaxy are illustrated in Figure 11 (top) as red points, with dusty source candidates signified with over-plotted green dots. Many of these objects are anticipated to be TP-AGB stars (other possibilities include massive YSOs, however positive classification will require SED model fitting; see §4.2.2 for more details), with the dustiest among them appearing to cluster around the large SF region located in the lower left (the SE lobe). Furthermore, these AGB star candidates show consistency with identifications from past studies utilizing HST (e.g., Östlin & Mouhcine, 2005).
@@
Stars identified as RSGs are a late evolutionary phase of higher-mass (10 M) MS progenitors which trace regions of recent star formation. These short-lived, cool, very luminous objects occupy color space coincident with TP-AGB stars, and are expected to be found alongside them in the F200W vs. F115W – F200W CMD’s evolved-star branch. The majority of the brightest sources populating the left-side branch (Figure 7; blue points), however, are denoted as being candidate dust-enshrouded stars. These are therefore likely RSG candidates as well, where they are illustrated in Figure 11 (bottom) as blue points with over-plotted green dots. We find that these dusty left-branch RSG candidates are strongly clustered around the large SF region located in the upper right (the NW lobe), opposite to that of the dusty TP-AGB stars discussed earlier. This broad demarcation of differently-aged sources suggests that the bulk of the most recent star-formation episode of the SE lobe, populated predominantly with slower-evolving C-rich AGB stars, took place longer ago than that of the NW lobe, which is still inhabited by younger sources such as RSGs.
@@ 4.2.2 Young Stellar Object Candidates
The active star formation taking place in I Zw 18 makes it an excellent local analog for conditions which pervaded the early universe, when heavy-element enrichment levels had yet to reach what is found today. YSOs are birthed in active star-formation regions and exhibit strong IR-excess as light is absorbed and re-emitted by cool, dusty envelopes and accretion disks (Meixner et al., 2013; Seale et al., 2014). Identification of YSOs in I Zw 18 represents an improvement in our understanding of star formation mechanisms at the very low levels of chemical enrichment reminiscent of the early universe, as resolved studies in higher-redshift galaxies that are comparably metal-poor are observationally precluded in all but a few rare circumstances. These YSOs trace regions of active star formation and emit even more strongly at long wavelengths than dusty AGB stars, particularly at the earliest stages of their lifetimes. This disk and its circumstellar envelope disperses as the protostar evolves (Robitaille et al., 2006), and so by consequence its dominant emission shifts from mid- to near-IR.
We identify YSO candidates in I Zw 18 as point sources exhibiting very red colors, similar to those of dust-enshrouded TP-AGB stars. This is accomplished by implementing the strategy featured in Jones et al. (2023) and Habel et al. (submitted) for JWST study of YSOs in the SMC SF region NGC 346. The selection criteria requires the source to be redward of the evolved star branch from the F200W vs. F115W – F200W CMD (with color F115W – F200W 1.2). Additionally, the YSO candidates must satisfy the condition of F200W – F444W 1.0, which was used for NGC 346 to establish the presence of IR excess. Finally, in order to correct for source contamination from background galaxies, which possess similar IR colors, we perform a final visual inspection of the YSO candidates in the NIRCam F115W image to reject extended objects. We present these results in the color-color diagram (CCD) as Figure 12 (left), with = 15 point sources plotted in green which inhabit the dashed line box illustrating these parameters. The location of these YSO candidates on the F200W vs. F115W – F200W primary diagnostic CMD (Figure 12; right) emphasizes the characteristic very red colors, as well as some magnitudes fainter than the TRGB (F200W = 25.6). We note that due to limitations of resolution and sensitivity at the distance to I Zw 18, any YSO candidates putatively identified in this study are likely to be unresolved clusters of several young stars.
@@ 4.3 Component C
The companion system to I Zw 18, Component C, is offset from the main galaxy by 36 arcsec. We present a comparison of the F200W vs. F115W – F200W CMD with Component C point sources (open red circles) overlaid upon stars found in the main body (gray dots) as Figure 14.
Contrasting the distributions of sources, we see stellar populations which are roughly self-consistent between I Zw 18 and Component C ( = 145), forming distinctive left- and right-side branches which represent young and evolved stars, respectively. Of particular note is a comparative lack of Component C sources populating the bright end of the left-side branch, where the youngest UMS stars are expected to be located. This suggests that Component C is home to a generally-older population of stars, where the youngest and hottest UMS sources have already transitioned to later evolutionary stages. An investigation of dusty targets, including selecting objects above the F200W vs. F115W – F200W CMD TRGB (F200W 25.6) and with F115W – F444W color 0.5, yields a small number ( = 26) of sources denoted as filled red circles in Figure 14. These appear to predominantly inhabit the bright end of the F200W vs. F115W – F200W CMD’s evolved star branch. See Bortolini et al. (in prep.) for an exploration of the SFH for Component C.

### v2

@@ 4.1 Images
Our NIRCam-only image (Figure 2) reveals an extreme population of bright, blue, recently formed massive stars, with positions centered upon two primary lobes of star formation, consistent with earlier optical/near-IR studies of I Zw 18 utilizing HST (e.g., Hunter & Thronson, 1995; Östlin, 2000; Cannon et al., 2002; Izotov & Thuan, 2004). An assortment of vivid red stellar sources distributed about the galaxy accentuates the presence of luminous evolved stars. Also apparent are the large supershells of dust and ionized gas carved out by previous star-formation and SNe activity. These align with extant maps of emission, including continuum-subtracted H and H from Cannon et al. (2002), which found significant concentrations of dust along such structures. Resolved background galaxies are apparent throughout the image spanning a variety of orientations and morphology. Stellar sources associated with the companion system Component C are visible at the top of the image, just left of center. Gas structure of the ISM is noticeably absent by comparison, consistent with a generally older population of stars bereft of recent star formation (see §4.3 for further analysis).
@@
Single-filter false-color images for each of the four NIRCam (Figure 4) and four MIRI (Figure 5) filters are presented which emphasize details observed in each band. The spatial distribution of all point sources extracted from both I Zw 18’s main body ( = 1,520; gray dots) and Component C ( = 286; red open circles) are shown in Figure 6.
4.2 Color–Magnitude Diagrams
@@
We construct near-IR CMDs using NIRCam data, which enabled our investigation of the various stellar populations inhabiting I Zw 18 (Figure 7). Our primary diagnostic CMD is F200W vs. F115W–F200W, which possesses the highest sensitivity to faint sources and therefore offers the largest potential pool of candidate stars ( = 689). This combination of filters was employed as a JWST-specific facsimile for standard diagnostic plots, enabling comparison with previous dusty and evolved star studies utilizing Spitzer bands such as SAGE and DUSTiNGS (e.g., Blum et al., 2006; Boyer et al., 2011, 2015a, 2015b). In addition, the methods executed by this study mirror recent stellar populations imaging work undertaken with JWST of other extragalactic sources, including NGC 6822 (Lenkić et al., 2024; Nally et al., 2024) and NGC 346 (Jones et al., 2023; Habel et al., 2024).
@@
The structure of points populating this CMD can be roughly described as two primarily vertical columns extending from a common foundation, which terminates as the detection threshold is reached at magnitudes fainter than F200W 28.0. At this level, the sensitivity limitations of the photometry results in a broad scatter. Points which make up the blueward (left-side) column of sources represent young stars of the upper main sequence (UMS), with those exhibiting the brightest magnitudes being the most massive. The redward column of sources is principally populated by evolved stars. Here, fainter sources are on the red giant branch (RGB), while at brighter luminosities these stars are chiefly He-core-burning RSGs and those on the AGB. Based on the distance modulus to I Zw 18, the tip of the red giant branch (TRGB) is estimated to be located at F200W 25.6. A luminosity function for all galaxy stars (gray) and only those stellar sources populating the evolved star column of sources in this CMD (red) is presented as Figure 8. An apparent discontinuity of sources at F200W 25.6 supports this estimation of the TRGB location, though a combination of large photometric uncertainties and small number statistics prevents definitive identification of the TRGB.
@@
In order to assist our preliminary point-source type classifications, we have included for reference parsec isochrones of an XMP stellar population ( = 0.0003; e.g., Bressan et al., 2012) at ages of 1, 10, 25, 50, 100, 250, 500 Myr, 1, 2, 5, and 10 Gyr, overlaid on this CMD (see Figure 9). The distribution of sources seen reflects the existence of both older and younger populations; for a thorough examination of the star formation history (SFH) of this galaxy as discerned by the photometric data obtained with NIRCam, see Bortolini et al. (2024). The overall numbers for point-source identifications described in the following subsections from the young star and evolved star branches, including their respective dust-enshrouded candidates (§4.2.1), plus YSO and YSO cluster candidates (§4.2.2), are summarized in Table 4.
@@ 4.2.1 Dusty Evolved Star Candidates
Identification of dust-enshrouded stellar source candidates in I Zw 18 using IR photometry was accomplished by examining a CMD employing our most widely separated NIRCam filter data (F444W vs. F115W–F444W; Figure 10). At these longer wavelengths, the numbers of detected point sources ( = 183) diminish in comparison to our primary diagnostic CMD. This large color baseline, however, affords a robust separation between dusty and nondusty stellar types: Any source detection with a color value F115W–F444W 0.0 may be considered to be at least somewhat enshrouded in dust. In order to account for photometric uncertainties, multiple ages, variance in metallicities, and other broadening effects, we have elected to implement a more conservative color cut at F115W–F444W = 0.5. Our dust-enshrouded source candidates therefore require detection in all three of the F115W, F200W, and F444W filters; a star becomes ineligible for inclusion in our high-confidence catalog if it is missing from any one filter’s detected source list. While the overall effect is that we are possibly missing some real dust-enshrouded source candidates in our study, we are opting to remain conservative. We also note that, while the MIRI F770W filter provides for a yet-longer color baseline, the paucity of sources detected via the F115W–F770W color for our data diminishes its utility as a diagnostic tool. At the distance to I Zw 18 (18.2 Mpc), the resolution and sensitivity of MIRI F770W are substantially less than that of NIRCam F444W, such that including this baseline could result in blending in the F770W filter, leading to source mismatches between catalogs.
The F444W vs. F115W–F444W CMD (Figure 10) employed to characterize dust-enshrouded source candidates in I Zw 18 bears resemblance to the [4.5] vs. –[4.5] CMD used by the Dell’Agli et al. (2019a) study of Sextans A, a nearby ( = 1.32 Mpc; Dolphin et al., 2003) low-metallicity (12 + (O/H) 7.6; Skillman et al., 1989; Pilyugin, 2001) SF dwarf galaxy (see their Figures 2 and 3). While Dell’Agli et al. (2019a) utilized this plot alongside evolutionary model tracks of metal-poor stars in order to make distinctions between evolved star populations (RSGs, O-rich AGBs, C-rich AGBs), we do not attempt to make equivalent early distinctions. Because of the larger distance to I Zw 18, our data suffers from higher photometric uncertainties as a consequence of crowding effects.
@@
For our list of dust-enshrouded source candidates identified from this CMD, we additionally required each detection to have been found in the F200W vs. F115W–F200W CMD located above the TRGB (F200W = 25.6). We note that, while some additional bona fide dust-enshrouded source candidates are likely to lie below this TRGB, we have elected to exclude them here for the sake of remaining conservative with our estimates, and to reflect larger photometric uncertainties at these magnitudes. For objects inhabiting the right-side (evolved stars) branch of the F200W vs. F115W–F200W CMD, those which are additionally identified as being dusty in the F444W vs. F115W–F444W CMD are expected to be AGB stars, RSGs, and bright YSO clusters. The fewer dusty sources of the left-side (young stars) branch of the F200W vs. F115W–F200W CMD are likely RSGs. In total, we find = 119 dusty source candidates in the F444W vs. F115W–F444W CMD, which are signified both in Figures 7 and 10 as overlaid green dots. Of these, = 99 belong to the evolved stars branch, while = 20 belong to the young stars branch. Overall, of the = 226 sources populating the evolved stars branch, the = 99 robust dust-enshrouded source candidates represent 43.8% of the total. While the = 20 dust-enshrouded source candidates on the young stars branch constitute only 17.2% of the total ( = 116), these include the majority of bright sources expected to be RSGs.
@@
Stars inhabiting the AGB represent the final evolutionary stage of low- to intermediate-mass (0.6–10 ) MS progenitors, and are believed to be a major contributor of interstellar dust. Of interest to this observing program in particular are thermally pulsing (TP-AGB) stars which lie above the TRGB. We may classify these AGB stars into two major categories: oxygen- and carbon-rich (O-rich and C-rich, respectively). MS progenitors with masses 3–8 evolve to become AGB stars which are rich in oxygen, but have short lifespans ( 30–200 Myr). While they are capable of producing dust relatively quickly after having formed, they are not expected to enrich the ISM with s-process elements in metal-poor environments to the same degree as AGB stars with lower initial masses (1–3 ), which are more prevalent (e.g., Lugaro et al., 2003; Karakas & Lugaro, 2016; Karakas, 2010). Fewer s-process elements may result in fewer nucleation sites for dust growth in the ISM. These stars possess longer lifespans ( 200 Myr–3.6 Gyr; Karakas & Lattanzio, 2003; Ventura et al., 2013; Boyer et al., 2015a), during which they experience repeated third dredge-up episodes (Kwok, 2000) that lead to their becoming increasingly carbon-enriched (Karakas & Lattanzio, 2014). O- and C-rich TP-AGB stars may therefore be distinguished from one another based on the dominant chemistries of their photospheres. If the value of the ratio of carbon to oxygen (C/O) is 1, the AGB star is C-rich, while a C/O value 1 is O-rich. Dust species born from these populations depend upon the dominant element, such that carbonaceous dust grains are produced by C-rich AGB stars, while silicates are produced by O-rich AGB stars.
The stellar mass value at which the distinction between O- and C-rich AGB stars occurs exhibits a dependence on metal abundance, decreasing from 3 as metallicity gets lower: The condition of C/O 1 is reached more easily for stars in environments with lower abundances of oxygen. In the case of extremely low metallicities, such as that exhibited by I Zw 18, the threshold for producing more C-rich stars may occur as low as 2.5 (Dell’Agli et al., 2018).
@@
Without the wider color baselines afforded by MIRI photometry, robust segregation between O- and C-rich AGB stars is not possible. Many of the evolved star branch sources identified in the F200W vs. F115W–F200W CMD (Figure 7; red points), however, are expected to be TP-AGB stars. In addition, considering the extremely low metallicity of I Zw 18, many of these TP-AGB candidate stars are expected to be C-rich. The distribution of dust-enshrouded evolved stars (red points with overplotted green dots) permeates the parameter space occupied by the F200W vs. F115W–F200W CMD’s evolved star branch, suggesting the presence of many dusty C-rich AGB star candidates in I Zw 18.
@@
The locations of the evolved star branch sources in this galaxy are illustrated in Figure 11 (top) as red points, with dusty source candidates signified with overplotted green dots. Many of these objects are anticipated to be TP-AGB stars (other possibilities include massive YSOs, however positive classification will require spectral energy distribution model fitting; see §4.2.2 for more details), with the dustiest among them appearing to cluster around the large SF region located in the lower left (the SE lobe). Furthermore, these AGB star candidates show consistency with identifications from past studies utilizing HST (e.g., Östlin & Mouhcine, 2005).
@@
Stars identified as RSGs are a late evolutionary phase of higher-mass (10 ) MS progenitors which trace regions of recent star formation. These short-lived, cool, very luminous objects occupy color space coincident with TP-AGB stars, and are expected to be found alongside them in the F200W vs. F115W–F200W CMD’s evolved star branch. The majority of the brightest sources populating the left-side branch (Figure 7; blue points), however, are denoted as being candidate dust-enshrouded stars. These are therefore likely RSG candidates, as well, where they are illustrated in Figure 11 (bottom) as blue points with overplotted green dots. We find that these dusty left-branch RSG candidates are strongly clustered around the large SF region located in the upper right (the NW lobe), opposite to that of the dusty TP-AGB stars discussed earlier. This broad demarcation of differently aged sources suggests that the bulk of the most recent star-formation episode of the SE lobe, populated predominantly with slower-evolving C-rich AGB stars, took place longer ago than that of the NW lobe, which is still inhabited by younger sources such as RSGs.
@@ 4.2.2 Young Stellar Object Candidates
The active star formation taking place in I Zw 18 makes it an excellent local analog for conditions which pervaded the early Universe, when heavy-element enrichment levels had yet to reach what is found today. YSOs are birthed in active star formation regions and exhibit strong IR excess as light is absorbed and reemitted by cool, dusty envelopes and accretion disks (Meixner et al., 2013; Seale et al., 2014). Identification of YSOs in I Zw 18 represents an improvement in our understanding of star formation mechanisms at the very low levels of chemical enrichment reminiscent of the early Universe, as resolved studies in higher-redshift galaxies that are comparably metal-poor are observationally precluded in all but a few rare circumstances. These YSOs trace regions of active star formation and emit even more strongly at long wavelengths than dusty AGB stars, particularly at the earliest stages of their lifetimes. This disk and its circumstellar envelope disperses as the protostar evolves (Robitaille et al., 2006), and so by consequence its dominant emission shifts from mid- to near-IR.
We identify YSO candidates in I Zw 18 as point sources exhibiting very red colors, similar to those of dust-enshrouded TP-AGB stars. This is accomplished by implementing the strategy featured in Jones et al. (2023) and Habel et al. (2024) for JWST study of YSOs in the SMC SF region NGC 346. The selection criteria require the source to be redward of the evolved stars branch from the F200W vs. F115W–F200W CMD (with color F115W–F200W 1.2). Additionally, the YSO candidates must satisfy the condition of F200W–F444W 1.0, which was used for NGC 346 to establish the presence of IR excess. Finally, in order to correct for source contamination from background galaxies, which possess similar IR colors, we perform a final visual inspection of the YSO candidates in the NIRCam F115W image to reject extended objects. We present these results in the color-color diagram (CCD) as Figure 12 (left), with = 15 point sources plotted in green which inhabit the dashed line box illustrating these parameters. The location of these YSO candidates on the F200W vs. F115W–F200W primary diagnostic CMD (Figure 12; right) emphasizes the characteristic very red colors, as well as some magnitudes fainter than the TRGB (F200W = 25.6). We note that due to limitations of resolution and sensitivity at the distance to I Zw 18, any YSO candidates putatively identified in this study are likely to be unresolved clusters of several young stars.
@@ 4.3 Component C
The companion system to I Zw 18, Component C, is offset from the main galaxy by 36″. We present a comparison of the F200W vs. F115W–F200W CMD with Component C point sources (open red circles) overlaid upon stars found in the main body (gray dots) as Figure 14.
Contrasting the distributions of sources, we see stellar populations which are roughly self-consistent between I Zw 18 and Component C ( = 145), forming distinctive left- and right-side branches that represent young and evolved stars, respectively. Of particular note is a comparative lack of Component C sources populating the bright end of the left-side branch, where the youngest UMS stars are expected to be located. This suggests that Component C is home to a generally older population of stars, where the youngest and hottest UMS sources have already transitioned to later evolutionary stages. An investigation of dusty targets, including selecting objects above the F200W vs. F115W–F200W CMD TRGB (F200W 25.6) and with F115W–F444W color 0.5, yields a small number ( = 26) of sources denoted as filled red circles in Figure 14. These appear to predominantly inhabit the bright end of the F200W vs. F115W–F200W CMD’s evolved stars branch. See Bortolini et al. (2024) for an exploration of the SFH for Component C.